OpenStudy (anonymous):

solve for x 8 -3x = 20 A. x=4 B x= 9.33 C x=-4 D x= -9.33

OpenStudy (anonymous):

8 - 3x = 20 -3x = 12 x = 12/-3 = -4 Option C.